William opened two investment accounts for his grandson's college fund. The first year, these investments, which totaled $18,000 yielded $831 in simple interest. Part of the money was invested at 5.5% and the rest at 4%. How much was invested at each rate?
Answer by mananth(16949)

part I 5.50% per annum---x
part II 4.00% per annum---y

x+y=18000 ------------------------1
5.50%x +4.00%y=831
Multiply by 100
5.5x+4y=83100.00--------2

Multiply (1) by -5.5
we get

-5.5 x -5.5 y= -99000.00

Add this to (2)

-1.5 y=-15900

divide by -1.5

y=10600
$10,600 investment at 4.00%
Balance $ 7400 investment at 5.50%
CHECK
10600 --------- 4.00% ------- 424.00
7400 ------- 5.50% ------- 407.00
Total -------------- 831.00